Sunday 25 March 2018

Maintenance Update (vacuum, plug, Hall, niggles)

Tuesday 20th
Warming fire required in readiness for a test run on Wednesday.

Wednesday 21st
I’m not sure who lit the steam-raising fire this morning - it could have been Jamie, Chris-1, Chris-2 or even John H. Anyway, there was 60 psi on the clock when I arrived at 10.30. Bruce and Gil were going round all of the oiling-up points. As the loco had been stood idle for months, all of the pots were empty. Hence it took all morning to complete this task.

Note that it takes three people to oil the vacuum pump - one to hold the oil can; one to tip it up, and one to manage the process. šŸ˜‰

Then disaster!

Plug no.18 was leaking. As pressure rose to 130 psi, there was a major plume emitting from the lower part of the plug. John H declared it a failure, and immediately began to deaden the fire to prevent pressure rising any more.

I find this extremely odd, as 2807 passed her steam test on 2nd March; Mark Y nipped up plugs that were wisping steam on 6th; and today this one is gushing steam dangerously. Has someone removed it and not fitted it back tightly while we were not looking, or what?

While there was pressure, JC ran her up & down a bit to check that there were no complaints from bearings, valves or cylinders. This all went very well - no problems at all.

The problem then was how to get pressure back down! 2807 is well known for holding the heat - maybe she has an efficient firebox/boiler. So, injecting (cold) water helped a little; opening the steam heating and venting it helped … a little; eventually a blow-down had to be undertaken to make room for more water in the boiler! By 4 pm, pressure was down at 15 psi.

Mark came in and replace plug 18 with a new one on Thursday. The hole’s thread needed re-cutting.

There were a few niggles that turned up. The slacking pipe (aka “pep” pipe) dribbles constantly. There was a drip from the condensing coil in the cab roof. The drain cock on the gauge frame doesn’t line up correctly. One of the glasses in the hydrostatic lubricator totally clogged up with oil.

The cock on the pep pipe is going to be replaced with an improved design within a few weeks, so we can ignore that.

Bruce fixed the condensing coil.

Bruce also sucked the oil out of the glass and cleaned that.

John G was here today, too, but it was difficult to know what to do, since it was “all go” and then “all stop!” However, he applied some elbow grease to the coupling rods, cleaning off some rust spots and places where paint had crept round from their backsides.

Apart from playing fireman for a while, I just needle-gunned four chairs.

Saturday 24th
Alex lit a warming fire on Friday, and Ade + Eleanor were crew today. Alex and Gwendolyn did a spot of cleaning (inside the cab had been in an awful state). The glasses on the hydrostatic lubricator kept oiling-up. JC replaced them - apparently some new glasses were not quite to spec.

In due course, steam was raised; the plug behaved, so she went for trip double-heading Foremarke Hall. Gil went along for the ride.

There were a couple of niggles (as usual) - she wouldn’t raise enough vacuum in the reservoir (so Foremarke helped). Bruce tightened the gland on the loco vacuum cylinder, so hopefully that will be OK now. The new rod bearings are working fine - not over-heating or anything.

Here she is heading back into Toddington (Foremarke is at the rear). Ade commented that he hardly needed any more regulator than usual even though there was an extra 125 tons at the back!

Apart from the loco, John T, Rob and I prepared some railwayana for a customer who is building a section of railway in his garden. Rob did a spot of welding, including some for Dinmore Manor. Then we turn our attention to boot scraper production, ending the day with four black bottoms, four angle-ground and one waiting its turn.

Issues logged and outstanding are:
5 Pep pipe leaking, Valve needs seating. {We’re awaiting a new type of valve}
6 Driver side, under cab side, main steam pipe to injector leaking at joint/sleeve.
8 Gauge frame drain-down tap out of line (vertical) when closed.
9 Oil feed pipe to crown on right hand side driver loose. Unable to tighten nut further.
Ditto driver’s side intermediate.
10 Both clack valves passing-by steam.

For those of you who would like to see 2807 in action, she is currently rostered for:
Train 3 on March 30 & 31
April 1 & 2; and then from 19 to 26 inclusive, plus 28 & 29.


No comments: